# A Generalized Sylvester-Fermat-Torricelli problem with application in disaster relief operations by UAVs

### v1

@@ 5. $\bullet$
which implies $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq N(\bar{x};\Omega_{r})\cap\Upsilon$.
For the opposite inclusion, take any $v\in N(\bar{x};\Omega_{r})\cap\Upsilon$, we need to show that the following inequality holds:
@@
By applying Proposition 3.4, we know that $v\in\partial\rho_{F}^{\Omega_{r}}(\bar{x})$, which gives:
@@
If $x\notin\Omega_{r}$, we can use Lemma 8.6, and then:
@@
which implies $v\in\partial\rho_{F}^{\Omega}(\bar{x})$.
In the other case, suppose $x\in\Omega_{r}$ or $\rho_{F}^{\Omega}(x)=p\leq r$. Since $\sigma_{F}(v)=1$, for any $\epsilon>0$, we can find $f\in F$ such that $\langle v,f\rangle\geq 1-\epsilon$. Using Lemma 8.7, we obtain $y\in\Omega_{r}$, where y=x+(rp)fy=x+(r-p)f. Furthermore, since $v\in N(\bar{x};\Omega_{r})$, it follows that $\langle v,y-\bar{x}\rangle\leq 0$. Therefore:
@@
$\Rightarrow\langle v,x-\bar{x}\rangle\leq(p-r)\langle v,f\rangle\leq(1-\epsilon)(p-r)=(1-\epsilon)(\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})).$
Hence, $v\in\partial\rho_{F}^{\Omega}(\bar{x})$ as $\epsilon\rightarrow 0$, and we obtain $N(\bar{x};\Omega_{r})\cap\Upsilon\subseteq\partial\rho_{F}^{\Omega}(\bar{x})$. ∎
Now, we need an important definition for the following proposition. Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. Then, the generalized projection from a point $x\in\mathbb{R}^{q}$ to the set $\Omega$, denoted by $\Pi_{F}(x;\Omega)$, is defined as follows:
$\Pi_{F}(x;\Omega):=\{\omega\in\Omega\ |\ \rho_{F}^{\Omega}(x)=\rho_{F}(x-w)\}.$ (3.5)
The nonemptiness of $\Pi_{F}(x;\Omega)$ is established in Proposition 8.1.
The following proposition provides an upper estimate for the subdifferential of the function $\rho_{F}^{\Omega}$, which is useful for computing the subdifferential at points where the function is differentiable.
Proposition 3.6 (Subdifferential estimation for $\rho_{F}^{\Omega}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int }F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty closed convex set. The subdifferential of the set-based Minkowski gauge function at $\bar{x}\in\textnormal{dom}\rho_{F}^{\Omega}$ can be estimated as follows:
$\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega}),$
where $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$. If, in addition, $\bar{x}\in\textnormal{int(dom}\rho_{F}^{\Omega})$ and $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$, then the inclusion becomes an equality:
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})=\{\nabla\rho_{F}(\bar{x}-\bar{\omega})\}.$
@@
Let $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. Then, by Proposition 3.1, it follows that:
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega)\leq\rho_{F}(x-\bar{\omega})-\rho_{F}(\bar{x}-\bar{\omega}).$
Therefore $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})$ and $\partial\rho_{F}^{\Omega}(\bar{x})\subseteq\partial\rho_{F}(\bar{x}-\bar{\omega})$.
For the final part of the proposition, we know that $\rho_{F}^{\Omega}$ is convex due to Proposition 3.2. Therefore, by Theorem 23.4 in [37], its subdifferential is nonempty at $\bar{x}\in\operatorname{int}(\operatorname{dom}\rho_{F}^{\Omega})$. Furthermore, since $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$, the subdifferential $\partial\rho_{F}(\bar{x}-\bar{\omega})$ reduces to a singleton (see Theorem 5.2 in [22]). It then follows that
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega}).$
Finally, by Theorem 25.1 in [37], we have $\partial\rho_{F}(\bar{x}-\bar{\omega})=\{\nabla\rho_{F}(\bar{x}-\bar{\omega})\}.$
The next proposition establishes the relationship between $\rho_{F}^{\Omega}$ and its positive scalar multiples, demonstrating how positive coefficients and weights can be incorporated into the dynamic set.
Proposition 3.7 (Relationship between $\rho_{F}^{\Omega}$ and its positive scalar multiple).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed convex set, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty set. Then, for $\lambda>0$ we have:
$\lambda\rho^{\Omega}_{F}(x)=\rho^{\Omega}_{\frac{F}{\lambda}}(x).$
@@
First, consider the case where $\rho_{F}^{\Omega}(x)<\infty$. Given $\epsilon>0$, choose $t$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\epsilon$. Then, for any $\epsilon$, there exist $\omega\in\Omega$ and $f\in F$ such that $x=\omega+tf$. We can rewrite this as $x=\omega+\lambda t\cdot\frac{f}{\lambda},$ which implies that $x\in\Omega+(\lambda t)\frac{F}{\lambda}$. Therefore, $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda t<\lambda(\rho_{F}^{\Omega}(x)+\epsilon).$ Letting $\epsilon\to 0$ yields $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda\rho_{F}^{\Omega}(x).$ For the reverse inequality, let $\alpha>0$, and choose $d$ such that $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq d<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha$. Then, for any $\alpha$, there exist $\omega\in\Omega$ and $\frac{f}{\lambda}\in\frac{F}{\lambda}$ such that $x=\omega+d\cdot\frac{f}{\lambda}.$ This implies $x\in\Omega+\frac{d}{\lambda}F$, and therefore, $\rho_{F}^{\Omega}(x)\leq\frac{d}{\lambda}<(\frac{1}{\lambda})(\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha)$. Multiplying through by $\lambda$ gives $\lambda\rho_{F}^{\Omega}(x)<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha$. Letting $\alpha\to 0$ yields $\lambda\rho_{F}^{\Omega}(x)\leq\rho_{\frac{F}{\lambda}}^{\Omega}(x)$. Combining both inequalities, we conclude that $\rho_{\frac{F}{\lambda}}^{\Omega}(x)=\lambda\rho_{F}^{\Omega}(x)$.
Next, consider the case where $\rho_{F}^{\Omega}(x)=\infty$, therefore, $\{t\geq 0\ |\ x\in\Omega+tF\}=\emptyset$. This implies that the set $\{t\geq 0\ |\ x\in\Omega+t\frac{F}{\lambda}\}=\emptyset$, so
$\lambda\rho_{F}^{\Omega}(x)=\infty=\rho_{\frac{F}{\lambda}}^{\Omega}(x).$
In the final part of this section, we define a new function that will be useful in practical applications.
Let $F\subset\mathbb{R}^{n}$ be a nonempty closed bounded convex set, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty bounded set; then, for a point $x\in\mathbb{R}^{q}$, the Maximal Set-based Minkowski Gauge (MSMG) function is defined as:
$\bar{\rho}_{F}^{\Omega}(x):=\inf\{t\geq 0\ |\ x-\Omega\subseteq tF\}.$ (3.6)
In simple terms, the set $x-\Omega$ includes all vectors that start from a point in $\Omega$ and end at $x$. If we view $F$ as a dynamic set, then $\bar{\rho}_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ from the farthest point in $\Omega$. This function has been investigated in the literature, where a slightly different notation, $C_{\Omega}^{F}(x)$, as defined in (8.8), is commonly employed. Lemma 8.14 formalizes the connection between $C_{\Omega}^{F}(x)$ and the MSMG function. The application of the MSMG function will be presented in Section 6. The following proposition demonstrates the relationship between the MSMG function and $\rho_{F}$.
Proposition 3.8 (Relationship between $\bar{\rho}_{F}^{\Omega}$ and $\rho_{F}$).
Let $F\subset\mathbb{R}^{q}$ be a nonempty closed bounded convex set with $0\in\textnormal{int }F$. Then, for a nonempty bounded set $\Omega\subset\mathbb{R}^{q}$, we have:
@@
First, consider the case where $\bar{\rho}_{F}^{\Omega}(x)=0$. This implies that $\Omega=\{x\}$. Consequently, $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)=0$, and the equality in (3.7) holds.
Next, consider the case $\bar{\rho}_{F}^{\Omega}(x)>0$, and let $\epsilon>0$ be arbitrary. Set $t=\bar{\rho}_{F}^{\Omega}(x)-\epsilon$, so that $t>0$. By the definition of $\bar{\rho}_{F}^{\Omega}(x)$, there exists $\omega\in\Omega$ such that $x-\omega\notin tF$. This implies that $t<\rho_{F}(x-\omega)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)$. Since $t\to\bar{\rho}_{F}^{\Omega}(x)$ as $\epsilon\to 0$, it follows that $\bar{\rho}_{F}^{\Omega}(x)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)$.
For the reverse inequality, let $d$ be such that $\bar{\rho}_{F}^{\Omega}(x)\leq d<\bar{\rho}_{F}^{\Omega}(x)+\alpha$, where $\alpha>0$ is arbitrary. By the definition of $\bar{\rho}_{F}^{\Omega}(x)$, this implies that $x-\omega\in dF$ for all $\omega\in\Omega$. Hence $\rho_{F}(x-\omega)\leq d$ for all $\omega\in\Omega$. Therefore $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq d<\bar{\rho}_{F}^{\Omega}(x)+\alpha$. As $\alpha\to 0$, $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\bar{\rho}_{F}^{\Omega}(x)$.
The following theorem outlines key properties of the MSMG function that are fundamental to its practical implementation.
Theorem 3.9 (Properties of the MSMG function).
Let $F\subset\mathbb{R}^{q}$ be a nonempty compact convex set with $0\in\textnormal{int}F$, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty compact convex set. For the MSMG function, we have:
  • i)
    The MSMG function is convex.
  • ii)
    The MSMG function is Lipschitz continuous.
  • iii)
    The subdifferential of the MSMG function can be estimated at $\bar{x}\in\textnormal{dom}\bar{\rho}_{F}^{\Omega}$ as follows:
    $\partial\rho_{F}(\bar{x}-\omega^{\prime})\subseteq\partial\bar{\rho}_{F}^{\Omega}(\bar{x}),$
    where $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega):=\{\omega\in\mathbb{R}\ |\ \bar{\rho}_{F}^{\Omega}(\bar{x})=\rho_{F}(\bar{x}-\omega)\}$.
  • iv)
    The subdifferential of the MSMG function can be obtained at $\bar{x}\in\textnormal{dom }\bar{\rho}_{F}^{\Omega}$ as follows:
    $\partial\bar{\rho}_{F}^{\Omega}(\bar{x})=\textnormal{co}\left\{\bigcup_{\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)}N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}\right\},$
    where $\Upsilon^{\prime}=\{v\in\mathbb{R}^{q}\ |\ \sigma_{-F}(v)=1\}$, $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)$ as defined in $(iii)$, and $r=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})$.
@@
For $(i)$, using Proposition 3.2, we obtain that $\rho_{F}$ is convex. By applying Propositions 3.8 and 2.1, we conclude that $\bar{\rho}_{F}^{\Omega}(\cdot)$ is convex. For $(ii)$, it follows from the subadditivity of the function $\rho_{F}$ (Theorem 6.14 in [22] provides its proof) we have:
$\Rightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\sup_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Rightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)-\sup_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Rightarrow\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
Also, for the function $\rho_{F}$, there exists a constant $l\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$, as shown in Proposition 6.18 of [22]. Then, we have:
$\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l(x-y).$
Therefore, $\bar{\rho}_{F}^{\Omega}(\cdot)$ is Lipschitz continuous due to the arbitrary choice of $x$ and $y$.
For (iiiiii), given the continuity of $\rho_{F}$ and the compactness of $\Omega$, we obtain $\bar{\Pi}_{F}(\bar{x};\Omega)\neq\emptyset$ by applying the Weierstrass Theorem. Take any $v\in\partial\rho_{F}(\bar{x}-\omega^{\prime})$; then we have:
@@
Therefore, $v\in\partial\bar{\rho}_{F}^{\Omega}(\bar{x})$, which implies that $\partial\rho_{F}(\bar{x}-\omega^{\prime})\subseteq\partial\bar{\rho}_{F}^{\Omega}(\bar{x}).$
To prove statement $(iv)$, we first show that
$\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime}),$
where $\omega^{\prime}:=\bar{\Pi}_{F}(\bar{x};\Omega)$. To this end, we apply Proposition 3.1 for $\rho_{-F}^{\{\bar{x}\}}(\omega)$, which gives $\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x}).$ Using Lemma 8.12, we have $\rho_{-F}(\omega-\bar{x})=\rho_{F}(\bar{x}-\omega),$. By the definition of $\bar{\Pi}_{F}(\bar{x};\Omega)$ and Proposition 3.8, we know that $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)$ satisfies
$\rho_{F}(\bar{x}-\omega^{\prime})=\sup_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega).$
Equivalently,
$\rho_{-F}(\omega^{\prime}-\bar{x})=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x}),$
and hence
@@
Using Proposition 8.10 and the equality in (3.8), we have:
@@
Then, by Proposition 3.5, we can write:
$\partial\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})=N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime},$
Finally, by applying Proposition 2.3, we obtain:
$\partial\bar{\rho}_{F}^{\Omega}(\bar{x})=\textnormal{co}\left\{\bigcup_{\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)}N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}\right\}.$
4 Existence and Uniqueness of Optimal Solutions in the Generalized SFT Problem and Its Weighted Variant
At the beginning of this section, we define a preliminary term. Given two points $x,y\in\mathbb{R}^{q}$, the line passing through them is defined by

### v2

@@ 5. $\bullet$
which justifies the inclusion $\subset$” in (3.5).
To verify the opposite inclusion in (3.5), pick any vector $v\in N(\bar{x};\Omega_{r})\cap\Upsilon$. We aim to check the fulfillment of the inequality
@@
Applying Theorem 3.4 tells us that $v\in\partial\rho_{F}^{\Omega_{r}}(\bar{x})$, which yields
@@
If $x\notin\Omega_{r}$, we use Lemma 8.5 from the Appendix and get
@@
which readily ensures that $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. In the other case, suppose that $x\in\Omega_{r}$, i.e., $\rho_{F}^{\Omega}(x)=p\leq r$. Fix any $\varepsilon>0$ and deduce from $\sigma_{F}(v)=1$ the existence of $f\in F$ such that $\langle v,f\rangle\geq 1-\varepsilon$. Using Lemma 8.6 from the Appendix, we obtain $y\in\Omega_{r}$ with y=x+(rp)fy=x+(r-p)f. Since $v\in N(\bar{x};\Omega_{r})$, it follows that $\langle v,y-\bar{x}\rangle\leq 0$. Therefore,
@@
$\Longrightarrow\langle v,x-\bar{x}\rangle\leq(p-r)\langle v,f\rangle\leq(1-\varepsilon)(p-r)=(1-\varepsilon)(\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})).$
Passing to the limit as $\varepsilon\downarrow 0$ gives us $v\in\partial\rho_{F}^{\Omega}(\bar{x})$. This shows that $N(\bar{x};\Omega_{r})\cap\Upsilon\subset\partial\rho_{F}^{\Omega}(\bar{x})$ and thus completes the proof of the theorem. ∎
In what follows, we need yet another evaluation of the subdifferential of the set-based Minkowski gauge function at out-of-set points. Given a compact and convex set $F\subset\mathbb{R}^{q}$ with $0\in\textnormal{int}\,F$, and given a nonempty, closed, and convex set $\Omega\subset\mathbb{R}^{q}$, the generalized projection from a point $x\in\mathbb{R}^{q}$ to the set $\Omega$ relative to $F$, denoted by $\Pi_{F}(x;\Omega)$, is defined by
$\Pi_{F}(x;\Omega):=\big\{\omega\in\Omega\ \big|\ \rho_{F}^{\Omega}(x)=\rho_{F}(x-w)\big\}.$ (3.7)
The nonemptiness of $\Pi_{F}(x;\Omega)$ is established in Proposition 8.1 from the Appendix.
Proposition 3.6.
Let $F$ and $\Omega$ be the sets from definition (3.7) of the generalized projection under the assumptions imposed therein, and let $\bar{x}\in{\rm dom}\,\rho^{\Omega}_{F}$. Then there is $\bar{\omega}\in\Pi_{F}(\bar{x};\Omega)$ with
$\partial\rho_{F}^{\Omega}(\bar{x})\subset\partial\rho_{F}(\bar{x}-\bar{\omega}).$ (3.8)
If in addition we have $\bar{x}\in\textnormal{int(dom}\,\rho_{F}^{\Omega})$ and the function $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$, then the inclusion in (3.8) becomes the equality
$\partial\rho_{F}^{\Omega}(\bar{x})=\partial\rho_{F}(\bar{x}-\bar{\omega})=\{\nabla\rho_{F}(\bar{x}-\bar{\omega})\}.$ (3.9)
@@
Picking $v\in\partial\rho_{F}^{\Omega}(\bar{x})$, we deduce from Proposition 3.1 that
$\langle v,x-\bar{x}\rangle\leq\rho_{F}^{\Omega}(x)-\rho_{F}^{\Omega}(\bar{x})=\inf_{\omega\in\Omega}\rho_{F}(x-\omega)-\inf_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega)\leq\rho_{F}(x-\bar{\omega})-\rho_{F}(\bar{x}-\bar{\omega})$
which gives us $v\in\partial\rho_{F}(\bar{x}-\bar{\omega})$ and thus verifies the inclusion in (3.8).
To prove the second part of the proposition, we deduce from the convexity of $\rho_{F}^{\Omega}$ by Proposition 3.2 and basic convex analysis that $\partial\rho_{F}^{\Omega}(\bar{x})\neq\emptyset$ if $\bar{x}\in\operatorname{int}(\operatorname{dom}\,\rho_{F}^{\Omega})$. Furthermore, the subdifferential $\partial\rho_{F}^{\Omega}(\bar{x}-\bar{\omega})$ reduces to a singleton if $\rho_{F}$ is differentiable at $\bar{x}-\bar{\omega}$. This readily leads us to (3.9) and thus completes the proof of all the claimed statements. ∎
Alternative calculations of subgradients of the set-based Minkowski gauge functions at out-of-state points, which involve the generalized projection (3.7) and the normal cone to the $r$-enlargement (3.2), are given in Theorem 8.10.
The next proposition establishes the precise relationship between $\rho_{F}^{\Omega}$ and its multiplication by positive scalars while demonstrating how positive coefficients and weights can be incorporated into the dynamic set.
Proposition 3.7.
Let $F\subset\mathbb{R}^{q}$ be a nonempty, closed, and convex set, and let $\Omega\subset\mathbb{R}^{q}$ be a nonempty set. Then we have the relationship
$\lambda\rho^{\Omega}_{F}(x)=\rho^{\Omega}_{\frac{F}{\lambda}}(x)\;\mbox{ for any }\;\lambda>0.$ (3.10)
@@
We begin with the case where $\rho_{F}^{\Omega}(x)<\infty$. Given $\varepsilon>0$, choose $t$ such that $\rho_{F}^{\Omega}(x)\leq t<\rho_{F}^{\Omega}(x)+\varepsilon$. Then there exist vectors $\omega\in\Omega$ and $f\in F$ with $x=\omega+tf$. Rewrite the latter in the form $x=\omega+\lambda t\cdot\frac{f}{\lambda}$ and get $x\in\Omega+(\lambda t)\frac{F}{\lambda}$, which implies in turn that $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda t<\lambda(\rho_{F}^{\Omega}(x)+\varepsilon)$. Letting $\varepsilon\downarrow 0$ gives us $\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\lambda\rho_{F}^{\Omega}(x)$.
To verify the reverse inequality in(3.10), pick $\alpha>0$ and choose a number $\beta$ such that
$\rho_{\frac{F}{\lambda}}^{\Omega}(x)\leq\beta<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha.$
Then there are $\omega\in\Omega$ and $\lambda^{-1}f\in\lambda^{-1}F$ with $x=\omega+\lambda^{-1}\beta f$. This yields $x\in\Omega+\lambda^{-1}\beta F$, which leads us to the inequalities
$\rho_{F}^{\Omega}(x)\leq\lambda^{-1}\beta<\lambda^{-1}\Big(\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha\Big),$
which yield $\lambda\rho_{F}^{\Omega}(x)<\rho_{\frac{F}{\lambda}}^{\Omega}(x)+\alpha$. Letting $\alpha\downarrow 0$ implies that $\lambda\rho_{F}^{\Omega}(x)\leq\rho_{\frac{F}{\lambda}}^{\Omega}(x)$ and thus justifies the claimed relationship in (3.10).
It remains to consider the case where $\rho_{F}^{\Omega}(x)=\infty$ and therefore $\{t\geq 0\ |\ x\in\Omega+tF\}=\emptyset$. The latter readily implies that $\{t\geq 0\ |\ x\in\Omega+t\frac{F}{\lambda}\}=\emptyset$ and thus
$\lambda\rho_{F}^{\Omega}(x)=\infty=\rho_{\frac{F}{\lambda}}^{\Omega}(x),$
which completes the proof of the proposition. ∎
To proceed further, we introduce a new function important in the subsequent applications. Let $F\subset\mathbb{R}^{n}$ be a nonempty, closed, bounded, and convex set, and let $\Omega\subset\mathbb{R}^{q}$ be nonempty and bounded. The maximal set-based Minkowski gauge (MSMG) is defined by
$\bar{\rho}_{F}^{\Omega}(x):=\inf\big\{t\geq 0\ \big|\ x-\Omega\subset tF\big\},\quad x\in\mathbb{R}^{q}.$ (3.11)
In simple terms, the set $x-\Omega$ includes all vectors that start from a point in $\Omega$ and end in $x$. If we view $F$ as a dynamic set, then $\bar{\rho}_{F}^{\Omega}(x)$ represents the minimal time required to reach $x$ from the farthest point in $\Omega$, see Fig. 7(a) as an example. Lemma 8.13 from the Appendix provides the connection between $\bar{\rho}_{F}^{\Omega}(x)$ and the maximal time function $C_{\Omega}^{F}(x)$ defined in (8.10), The application of the MSMG function will be presented in Section 6. Now we derive a relationship between (3.11) and the classical Minkowski gauge function $\rho_{F}$.
Proposition 3.8.
Let $F\subset\mathbb{R}^{q}$ be a compact and convex set with $0\in\textnormal{int }F$, and let $\Omega\subset\mathbb{R}^{q}$ be nonempty and bounded. Then we have the representation
@@
In the case where $\bar{\rho}_{F}^{\Omega}(x)=0$, we clearly have $\Omega=\{x\}$. Thus $\sup_{\omega\in\Omega}\rho_{F}(x-\omega)=0$, which justifies the equality (3.12) in this case. If $\bar{\rho}_{F}^{\Omega}(x)>0$, pick an arbitrary number $\varepsilon>0$ and put $t:=\bar{\rho}_{F}^{\Omega}(x)-\varepsilon>0$. It follows from the definition of $\bar{\rho}_{F}^{\Omega}(x)$ that there exists $\omega\in\Omega$ such that $x-\omega\notin tF$, which implies that
$t<\rho_{F}(x-\omega)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega).$
Since $t\to\bar{\rho}_{F}^{\Omega}(x)$ as $\varepsilon\downarrow 0$, we arrive at $\bar{\rho}_{F}^{\Omega}(x)\leq\sup_{\omega\in\Omega}\rho_{F}(x-\omega)$.
To verify the reverse inequality in (3.12), let $\beta$ be such that $\bar{\rho}_{F}^{\Omega}(x)\leq\beta<\bar{\rho}_{F}^{\Omega}(x)+\alpha$, where $\alpha>0$ is arbitrary. The definition of $\bar{\rho}_{F}^{\Omega}(x)$ tells us that $x-\omega\in\beta F$ for all $\omega\in\Omega$, and hence $\rho_{F}(x-\omega)\leq\beta$ on $\Omega$. Therefore, we get
$\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\beta<\bar{\rho}_{F}^{\Omega}(x)+\alpha,$
which yields the claimed reverse estimate in (3.12) as $\alpha\downarrow 0$ and thus completes the proof. ∎
The final theorem of this section reveals the key properties of the MSMG function (3.11) that are fundamental for our practical applications given below.
Theorem 3.9.
Let $F\subset\mathbb{R}^{q}$ be a compact and convex set with $0\in\textnormal{int}\,F$, and let $\Omega\subset\mathbb{R}^{q}$ be nonempty, compact, and convex. Then we have the following assertions:
(i) The MSMG function is convex.
(ii) The MSMG function is Lipschitz continuous on $\mathbb{R}^{q}$.
(iii) The subdifferential of the MSMG function at $\bar{x}\in\textnormal{dom}\bar{\rho}_{F}^{\Omega}$ satisfies the inclusion
$\partial\rho_{F}(\bar{x}-\omega^{\prime})\subset\partial\bar{\rho}_{F}^{\Omega}(\bar{x}),$
where $\omega^{\prime}\in\overline{\Pi}_{F}(\bar{x};\Omega):=\{\omega\in\mathbb{R}\ |\ \bar{\rho}_{F}^{\Omega}(\bar{x})=\rho_{F}(\bar{x}-\omega)\}$.
(iv) The subdifferential of the MSMG function at $\bar{x}\in\textnormal{dom}\,\bar{\rho}_{F}^{\Omega}$ is represented by
$\partial\bar{\rho}_{F}^{\Omega}(\bar{x})=\textnormal{co}\left\{\bigcup_{\omega^{\prime}\in\overline{\Pi}_{F}(\bar{x};\Omega)}N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}\right\},$ (3.13)
where $\Upsilon^{\prime}:=\{v\in\mathbb{R}^{q}\ |\ \sigma_{-F}(v)=1\}$, $\omega^{\prime}\in\overline{\Pi}_{F}(\bar{x};\Omega)$, and $r:=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})$.
@@
It follows from Proposition 3.2 that $\rho_{F}$ is convex. Then we deduce from Propositions 3.8 and 2.1 that $\bar{\rho}_{F}^{\Omega}(\cdot)$ is convex as well, which is claimed in (i). To verify (ii), we have by [mordukhovich2023easy, Theorem 6.14] that $\rho_{F}$ is subadditive. Combining this with Proposition 3.8 ensures the implications
$\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\rho_{F}(y-\omega)\quad\text{for all }\omega\in\Omega$
$\Longrightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)\leq\rho_{F}(x-y)+\sup_{\omega\in\Omega}\rho_{F}(y-\omega)$
$\Longrightarrow\sup_{\omega\in\Omega}\rho_{F}(x-\omega)-\sup_{\omega\in\Omega}\rho_{F}(y-\omega)\leq\rho_{F}(x-y)$
$\Longrightarrow\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y).$
It follows from [mordukhovich2023easy, Proposition 6.18] that there exists a constant $i\geq 0$ such that $\rho_{F}(x)\leq l\|x\|$. Therefore, we have
$\bar{\rho}_{F}^{\Omega}(x)-\bar{\rho}_{F}^{\Omega}(y)\leq\rho_{F}(x-y)\leq l(x-y),$
which justifies the Lipschitz continuity of $\bar{\rho}_{F}^{\Omega}(\cdot)$ on $\mathbb{R}^{q}$ due to the arbitrary choice of $x$ and $y$.
To verify assertion (iii), observe that $\overline{\Pi}_{F}(\bar{x};\Omega)\neq\emptyset$ as it follows from the classical Weierstrass existence theorem due to the continuity of $\rho_{F}$ and the imposed compactness of $\Omega$. Picking any subgradient $v\in\partial\rho_{F}(\bar{x}-\omega^{\prime})$, we have
@@
Therefore, $v\in\partial\bar{\rho}_{F}^{\Omega}(\bar{x})$, which implies that $\partial\rho_{F}(\bar{x}-\omega^{\prime})\subset\partial\bar{\rho}_{F}^{\Omega}(\bar{x})$ as claimed in (iii).
It remains to prove assertion (iv). Let us first show that
$\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime}),$ (3.14)
where $\omega^{\prime}\in\overline{\Pi}_{F}(\bar{x};\Omega)$. To this end, apply Proposition 3.1 to $\rho_{-F}^{\{\bar{x}\}}(\omega)$, which gives us $\sup_{\omega\in\Omega}\rho_{-F}^{\{\bar{x}\}}(\omega)=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x}).$ Using Lemma 8.11 from the Appendix, we get $\rho_{-F}(\omega-\bar{x})=\rho_{F}(\bar{x}-\omega)$. It follows from the definition of $\overline{\Pi}_{F}(\bar{x};\Omega)$ and Proposition 3.8 that any fixed generalized projection vector $\omega^{\prime}\in\bar{\Pi}_{F}(\bar{x};\Omega)$ satisfies the equality
$\rho_{F}(\bar{x}-\omega^{\prime})=\sup_{\omega\in\Omega}\rho_{F}(\bar{x}-\omega),$
which can be equivalently written as
$\rho_{-F}(\omega^{\prime}-\bar{x})=\sup_{\omega\in\Omega}\rho_{-F}(\omega-\bar{x})$
and tells us therefore that
@@
Employing Proposition 8.9 from the Appendix together with (3.15) leads us to the equalities
@@
Finally, it follows from Theorem 3.5 that
$\partial\rho_{-F}^{\{\bar{x}\}}(\omega^{\prime})=N(\omega^{\prime};\bar{x}-rF)\cap\Upsilon^{\prime}.$
This allows us to deduce from the subdifferential maximum rule in Theorem 2.3 the fulfillment of representation (3.13) and thus complete the proof of this theorem. ∎
4 Well-Posedness of the Generalized SFT Model
In this section, we verify the well-posedness of our main generalized SFT model and hence of its specification, meaning by well-posedness the existence and uniqueness of optimal solutions. Given two points $x,y\in\mathbb{R}^{q}$, denote the straight line passing through them by